OpenStudy (anonymous):

Use the quadratic formular to solve the equation 1/4 x^2 - 1/2 x=1

OpenStudy (anonymous):

(1/4)x^2-(1/2)x-1=0 substitute a=1/4 b=-1/2 c=-1 in formula: \[\LARGE x_{1/2}=\frac{ -b \pm\sqrt{b^2-4ac }}{2a}\]

OpenStudy (anonymous):

1/2(+)(-)under root 5/4 / (1/2)

OpenStudy (anonymous):

Is this correct? 1-sqrt5 , 1+sqrt5

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=solve+%281%2F4%29+x^2+-+%281%2F2%29+x%3D1 yes it's correct.

jhonyy9 (jhonyy9):

1/4 x^2 -1/2 x =1 x^2 -2x -4=0 x_1,2=(2+/- sqrt(4+16))/2 =(2+/- 2sqrt5)/2 = 1+/- sqrt5